OpenStudy (anonymous):

A ball is dropped from the top of a 1,000 foot building. The height of the ball is half its original height after each bounce. What will be the height of the ball after 12 bounces? I think it would be either geometric series or sequence. Not sure.

OpenStudy (campbell_st):

this is a geometric series where the 1st term is 1000 and the common ratio r = 1/2

OpenStudy (campbell_st):

so a term in a geometric series is \[a_{n} = a_{1} \times r^{n - 1}\] just substitute and evaluate to find the answer

OpenStudy (anonymous):

@campbell_st the formula I have is an=a1-a1 r^n/ 1-r